# What Ails Generative Structure-based Drug Design: Too Little or Too Much Expressivity?

### v1

@@ Appendix C Chemical diversity with fixed unlabelled graphs
We performed the following experiment. We randomly sampled 100 molecules from the ZINC250k dataset. For each molecule, we used its template unlabelled graph to generate novel valid molecules using the procedure described in Appendix A.3. Whenever we generated a duplicate molecule we tried again. We terminated the procedure when we generated 1000 unique molecules or reached the limit of 10 failed trials due to duplicate generation.
@@
Furthermore, we computed the average Tanimoto similarity between the original molecule and the newly generated ones sharing the unlabelled graph. We found that the average Tanimoto similarity was 0.28 ± 0.08 across the 100 sampled molecules from ZINC250k. We compared this number with the average pairwise Tanimoto similarity of the 100 molecules original molecules, which was 0.24 (within one standard deviation). We also contrast this number with the 0.85 similarity threshold, above which molecules have a high probability of having the same activity [Patterson et al., 1996].
@@ Appendix D Binding affinity approximation model details
In this section we provide more details of the scoring model from Eq. (4). Recall that we want to approximate the binding affinity estimation defined in Eq. (9) using the protein pocket information and ligand’s unlabelled graph structure. To that end we represent the protein pocket as a graph with residues as nodes. Each node is equipped with a one-hot encoding of a residue type (1 of 20 possible amino acids) as a feature vector and 3D coordinates computed as a center of mass of the corresponding residue. Two nodes are connected with an edge if they are at most 15Å apart and we limit each node to have at most 24 neighbours. We do not use any edge features. To compute the protein pocket embedding, we use E(3)-Equivariant Graph Neural Networks (EGNNs) [Satorras et al., 2021].
@@
where we use $L=3$ in our experiments. $\phi_{e},\phi_{x}$ and $\phi_{h}$ are all implemented as concatenation of the input followed by a fully connected 2-layer MLP with hidden and output dimensions of $16$ and the SiLU activation function [Hendrycks and Gimpel, 2016].
@@ D.1 Training details
The dataset for training the scoring model consists of triples $(G^{P},G^{M},y)$ representing the protein, ligand and the binding affinity estimated with Vina respectively. We randomly select 10% of the protein pockets for validation and the remaining 90% for training. The model is trained to minimize the mean squared error with the Adam optimizer [Kingma and Ba, 2014] with default parameters of its PyTorch implementation [Paszke et al., 2019]. The model was trained with batch size 128 for 10000 steps, after which we did not observe improvements in the validation loss. Model training took approximately 45 minutes on an Apple M1 CPU.
@@ Appendix F Ligand’s center of mass predictor
As mentioned in Section 4.3, in order to accurately estimate the binding affinity for a protein-ligand pair with Vina, the ligand’s 3D configuration must be in the vicinity of the protein 3D configuration. The reason for that is that the scoring function defined in Eq. (8) vanishes for atom pairs that are far apart [Trott and Olson, 2010]. Consequently, gradient based optimization used to find the best 3D configuration (Eq. (9)) will return the original configuration, because all gradients are zero whenever the protein and ligand are too far apart.
@@
where $\mathbf{x}^{(L)}_{v}$ is defined in Eq. (12). As shown in [Satorras et al., 2021], $\hat{\mathbf{x}}$ is then equivariant to translations and rotations as required for the center of mass predictor model. In our experiments, we used $L=4$ layers and a hidden dimension of $16$. The model has 8k trainable parameters and both its number of parameters and runtime were included when reporting the results for SimpleSBDD models in Table 1.
To train the model, we used the training protein-ligand pairs of the CrossDocked2020 [Francoeur et al., 2020] dataset described in Section 5. We further split them into train and validation sets to monitor overfitting using a 85-15 train-validation split. Similarly to [Peng et al., 2022], the split was performed to ensure that all validation protein pockets will have sequence similarity lower than 30% to all training protein pockets.
The model was trained with Adam optimizer [Kingma and Ba, 2014] with default parameters of its PyTorch implementation [Paszke et al., 2019] to minimize the Euclidean distance between the predicted and ground truth centers of mass. We used a batch size of $64$ and trained the model for 10000 steps until the validation loss converged to $\sim 1.15$Å. Training took approximately 25 minutes on an Apple M1 CPU.
@@ Appendix G Additional molecular metrics
Here we provide additional metrics to complement the ones reported in Table 1, which were omitted for presentation clarity. Namely, we include the Lipinski’s Rule of Five and LogP, which were defined in Section 2. For Lipinski’s Rule of Five, $5$ is the highest possible value. For LogP we do not indicate which method is the best as there is no straightforward criterion to compare LogP values. In [Peng et al., 2022] it is noted that values in the $(-0.4,5.6)$ interval are considered good drug candidates. Note that for SimpleSBDD, the value of LogP is the closest to the midpoint of that interval $(2.6)$. None of the other baseline methods reported LogP or the Lipinski’s Rule of Five.
@@ G.1 Novelty of generated molecules
Table 3: Additional metrics.
@@ Appendix I SimpleSBDD with property optimization
Table 4: Our method allows for simultaneous optimization of binding and other properties. Comparison of our method w/ and w/o property optimization. With explicit property optimization, we can significantly improve all metrics without sacrificing binding affinity. The increased sampling time is still an order of magnitude lower than baseline methods even though running exclusively on CPU.
@@
As mentioned in Section 5.2, our scoring model can be used for drug repurposing. We found that for each protein pocket it takes around 2 seconds to find 100 diverse molecules with much higher predicted binding affinity and other improved molecular metrics. Furthermore, we repeated the experiment with ChEMBL dataset [Mendez et al., 2018, Davies et al., 2015] and found that $g_{\theta}$ generalizes well across molecular databases. See Table 5 for more details
Table 5: Scoring model generalizes across datasets and quickly finds diverse high quality molecules. The scoring model $g_{\theta}$ used for drug repurposing. We score the molecular databases: ZINC250k and ChEMBL respectively with $g_{\theta}$. The best scoring ones are compared with the reference molecule.
@@
A useful construct for determining indistinguishability is that of computation trees [Jegelka, 2022]:
@@
Since the multiset of depth-$L$ computation trees represents all information LU-GNNs can extract [Jegelka, 2022], we see that

### v2

@@ Appendix C CHEMICAL DIVERSITY WITH FIXED UNLABELLED GRAPHS
We performed the following experiment. We randomly sampled 100 molecules from the ZINC250k dataset. For each molecule, we used its template unlabelled graph to generate novel valid molecules by randomly replacing atom types without violating valency constraints. Whenever we generated a duplicate molecule we tried again. We terminated the procedure when we generated 1000 unique molecules or reached the limit of 10 failed trials due to duplicate generation.
@@
Furthermore, we computed the average Tanimoto similarity between the original molecule and the newly generated ones sharing the unlabelled graph. We found that the average Tanimoto similarity was 0.28 ± 0.08 across the 100 sampled molecules from ZINC250k. We compared this number with the average pairwise Tanimoto similarity of the 100 molecules original molecules, which was 0.24 (within one standard deviation). We also contrast this number with the 0.85 similarity threshold, above which molecules have a high probability of having the same activity (Patterson et al.,, 1996).
@@ Appendix D BINDING AFFINITY APPROXIMATION MODEL DETAILS
In this section we provide more details of the scoring model from Eq. (4). Recall that we want to approximate the binding affinity estimation defined in Eq. (9) using the protein pocket information and ligand’s unlabelled graph structure. To that end we represent the protein pocket as a graph with residues as nodes. Each node is equipped with a one-hot encoding of a residue type (1 of 20 possible amino acids) as a feature vector and 3D coordinates computed as a center of mass of the corresponding residue. Two nodes are connected with an edge if they are at most 15Å apart and we limit each node to have at most 24 neighbours. We do not use any edge features. To compute the protein pocket embedding, we use E(3)-Equivariant Graph Neural Networks (EGNNs) (Satorras et al.,, 2021).
@@
where we use $L=3$ in our experiments. $\phi_{e},\phi_{x}$ and $\phi_{h}$ are all implemented as concatenation of the input followed by a fully connected 2-layer MLP with hidden and output dimensions of $16$ and the SiLU activation function (Hendrycks and Gimpel,, 2016).
@@ D.1 Training details
The dataset for training the scoring model consists of triples $(G^{P},G^{M},y)$ representing the protein, ligand and the binding affinity estimated with Vina respectively. We randomly select 10% of the protein pockets for validation and the remaining 90% for training. The model is trained to minimize the mean squared error with the Adam optimizer (Kingma and Ba,, 2014) with default parameters of its PyTorch implementation (Paszke et al.,, 2019). The model was trained with batch size 128 for 10000 steps, after which we did not observe improvements in the validation loss. Model training took approximately 45 minutes on an Apple M1 CPU.
@@ Appendix F LIGAND’S CENTER OF MASS PREDICTOR
As mentioned in Section 4.3, in order to accurately estimate the binding affinity for a protein-ligand pair with Vina, the ligand’s 3D configuration must be in the vicinity of the protein 3D configuration. The reason for that is that the scoring function defined in Eq. (8) vanishes for atom pairs that are far apart (Trott and Olson,, 2010). Consequently, gradient based optimization used to find the best 3D configuration (Eq. (9)) will return the original configuration, because all gradients are zero whenever the protein and ligand are too far apart.
@@
where $\mathbf{x}^{(L)}_{v}$ is defined in Eq. (10). As shown in (Satorras et al.,, 2021), $\hat{\mathbf{x}}$ is then equivariant to translations and rotations as required for the center of mass predictor model. In our experiments, we used $L=4$ layers and a hidden dimension of $16$. The model has 8k trainable parameters and both its number of parameters and runtime were included when reporting the results for SimpleSBDD models in Table 1.
To train the model, we used the training protein-ligand pairs of the CrossDocked2020 (Francoeur et al.,, 2020) dataset described in Section 5. We further split them into train and validation sets to monitor overfitting using a 85-15 train-validation split. Similarly to (Peng et al.,, 2022), the split was performed to ensure that all validation protein pockets will have sequence similarity lower than 30% to all training protein pockets.
The model was trained with Adam optimizer (Kingma and Ba,, 2014) with default parameters of its PyTorch implementation (Paszke et al.,, 2019) to minimize the Euclidean distance between the predicted and ground truth centers of mass. We used a batch size of $64$ and trained the model for 10000 steps until the validation loss converged to $\sim 1.15$Å. Training took approximately 25 minutes on an Apple M1 CPU.
@@ Appendix G ADDITIONAL MOLECULAR METRICS
Here we provide additional metrics to complement the ones reported in Table 1, which were omitted for presentation clarity. Namely, we include the Lipinski’s Rule of Five and LogP, which were defined in Section 2. For Lipinski’s Rule of Five, $5$ is the highest possible value. For LogP we do not indicate which method is the best as there is no straightforward criterion to compare LogP values. In (Peng et al.,, 2022) it is noted that values in the $(-0.4,5.6)$ interval are considered good drug candidates. Note that for SimpleSBDD, the value of LogP is the closest to the midpoint of that interval $(2.6)$. None of the other baseline methods reported LogP or the Lipinski’s Rule of Five.
@@ G.1 Novelty of generated molecules
Table 4: Additional metrics. Error bars correspond to the standard deviation across test protein pockets.
@@ Appendix I SIMPLESBDD WITH PROPERTY OPTIMIZATION
Table 5: Our method allows for simultaneous optimization of binding and other properties. Comparison of our method w/ and w/o property optimization. With explicit property optimization, we can significantly improve all metrics without sacrificing binding affinity. The increased sampling time is still an order of magnitude lower than baseline methods even though running exclusively on CPU. Error bars correspond to the standard deviation across test protein pockets.
@@
I.1 Ablation study
We analyzed how the number of proposal molecules $K$ impacts the performance. In Section 5.2 we chose $K=50$ and now investigate what happens for $K=10,20,50$. For this experiment, due to computational constraints, we considered a random subset of 20% of the test protein targets and generated 50 molecules per each (as opposed to the standard of 100). Note that the runtime of property optimization scales linearly with $K$.
We found that increasing $K$ improves both QED and SA as expected, as the model optimizes for a mixture of these two properties. Furthermore, we found that the choice of $K$ does not seem to affect the predicted binding affinity, which strengthens our point that the optimization of atom types for various properties for a fixed unlabeled graph structure does not degrade the predicted binding affinity. Please see Table 6 for more details.
Table 6: Increasing the number of proposals $K$ improves molecular metrics without degrading the predicted binding affinity.
Vina Score (↓) High Affinity (↑) QED (↑) SA (↑) Diversity (↑)
Test set -6.99 ± 2.16 - 0.48 ± 0.21 0.73 ± 0.14 -
SimpleSBDD–$\mathcal{PO}\ (K=10)$7.935 ± 1.469 0.705 ± 0.356 0.760 ± 0.107 0.715 ± 0.078 0.651 ± 0.054
SimpleSBDD–$\mathcal{PO}\ (K=20)$7.934 ± 1.465 0.724 ± 0.352 0.781 ± 0.100 0.724 ± 0.079 0.651 ± 0.055
SimpleSBDD–$\mathcal{PO}\ (K=50)$7.939 ± 1.462 0.715 ± 0.375 0.796 ± 0.094 0.734 ± 0.077 0.652 ± 0.055
@@ I.1 Ablation study
As mentioned in Section 5.2, our scoring model can be used for drug repurposing. We found that for each protein pocket it takes around 2 seconds to find 100 diverse molecules with much higher predicted binding affinity and other improved molecular metrics. Furthermore, we repeated the experiment with ChEMBL dataset (Mendez et al.,, 2018; Davies et al.,, 2015) and found that $g_{\theta}$ generalizes well across molecular databases. See Table 7 for more details
Table 7: Scoring model generalizes across datasets and quickly finds diverse high quality molecules. The scoring model $g_{\theta}$ used for drug repurposing. We score the molecular databases: ZINC250k and ChEMBL respectively with $g_{\theta}$. The best scoring ones are compared with the reference molecule. Error bars correspond to the standard deviation across test protein pockets.
@@
A useful construct for determining indistinguishability is that of computation trees (Jegelka,, 2022):
@@
Since the multiset of depth-$L$ computation trees represents all information LU-GNNs can extract (Jegelka,, 2022), we see that
@@
K.1 Persistent Homology
Our claims in Propostion 2 also hold when the LUGNN model is additionally equipped with persistent homology (PH) information, which makes it strictly more expressive (Immonen et al.,, 2023). This is a strict generalization, because PH computes global topological information such as the number of connected components or the number of independent cycles that LUGNNs cannot compute.
Specifically, we can show that if the two ligand graphs are indistinguishable by LUGNNs+PH then the complex graphs are also indistinguishable with LUGNNs+PH.
Proof
Graphs $G_{1}$ and $G_{2}$ are indistinguishable by LUGNN+PH if and only if: 1) $G_{1}$ and $G_{2}$ are indistinguishable by LUGNN, and 2) $G_{1}$ and $G_{2}$ have the same Betti numbers $\beta_{0}$, $\beta_{1}$, where
  • $\beta_{0}(G)$ is the number connected components of $G$, and
  • $\beta_{1}(G)=\beta_{0}(G)+|E|-|V|$, where $G=(V,E)$, i.e. $V$ is the set of vertices and $E$ is the set of edges.
We now assume that $G_{1}$ and $G_{2}$ are indistinguishable by LUGNNs and $\beta_{0}(G_{1})=\beta_{0}(G_{2})$ and $\beta_{1}(G_{1})=\beta_{1}(G_{2})$.
In proposition 2 we already showed that the complex graphs $C_{1}$ and $C_{2}$ are indistinguishable by LUGNNs, so it suffices to show that $C_{1}$ and $C_{2}$ have the same Betti numbers.
From our definition of the complex graph we see that it is always connected, i.e. has a single connected component, and thus $\beta_{0}(C_{1})=\beta_{0}(C_{2})=1$.
To evaluate the 1-st order Betti numbers we denote $G_{1}=(V_{1},E_{1})$, $G_{2}=(V_{2},E_{2})$, $C_{i}=\mathcal{C}(P,G_{i})$ for some arbitrary protein graph $P=(V_{P},E_{P})$ using the notation from the paper.
Since $G_{1}$ and $G_{2}$ are indistinguishable by LUGNNs+PH, we must have $|V_{1}|=|V_{2}|$ and $|E_{1}|=|E_{2}|$. Therefore
$\beta_{1}(C_{1})$ $=\beta_{0}(C_{1})+|E_{P}|+|E_{1}|+|E_{P}|\cdot|E_{1}|-|V_{P}|-|V_{1}|$
$=1+|E_{P}|+|E_{2}|+|E_{P}|\cdot|E_{2}|-|V_{P}|-|V_{2}|$
$=\beta_{1}(C_{2})$
and thus $C_{1}$ and $C_{2}$ are indistinguishable by LUGNNs+PH.
@@
Appendix M LICENSES
Datasets
  1. 1.
    CrossDocked2020 (Francoeur et al.,, 2020): GPL-2.0 license
  2. 2.
    ZINC250k (Irwin et al.,, 2012): GPL-3.0+ license
  3. 3.
    ChEMBL (Mendez et al.,, 2018; Davies et al.,, 2015): CC BY-SA 3.0
Pre-trained models
  1. 1.
    MoFlow (Zang and Wang,, 2020): MIT License
  2. 2.
    DiffSBDD (Schneuing et al.,, 2022): MIT License
  3. 3.
    Pocket2Mol (Peng et al.,, 2022): MIT License
  4. 4.
    TargetDiff (Guan et al., 2023a, ): MIT License
  5. 5.
    DecompDiff (Guan et al., 2023b, ): CC BY-NC 4.0
  6. 6.
    RGA (Fu et al.,, 2022): License status unclear
  7. 7.
    3D-MCTS (Du et al.,, 2023): License status unclear
  8. 8.
    Autogrow4 (Spiegel and Durrant,, 2020): Apache License Version 2.0
Chemical software
  1. 1.
    RDKit (Landrum et al.,, 2020): BSD 3-Clause License
  2. 2.
    Vina (Alhossary et al.,, 2015): Apache License Version 2.0
  3. 3.
    Gnina (McNutt et al.,, 2021): GPL-2.0 license